Write the equation of the line in slope intercept form
Function table
X : 1 3 6
Y: 8 14 13

Answer by person24(13)

X : 1 3 6
Y: 8 14 13
Okay we need to pick two points so we can find our slope..
I choose (1,8) and (3,14)
Now lets use the equation ysub2-ysub1/xsub2-xsub1
(14-8)/(3-1)
6/2
3
your slope (m) is 3
slope intercept for is y=mx+b
so we now have our m
we need b
y=3x+b
to find B
we can pick a random point they have given us
I pick (1,8) again
plug and chug
where x=1 and y=8
8=3(1)+b
see now we can find for b since we only have one variable in that problem
8=3+b
subtract 3
5=b
now we can make our equationnn
y=3x+5 :D